OpenStudy (anonymous):

value of tan(-945)degree

OpenStudy (anonymous):

tan(-945)=tan(-945+720)=tan(-225)

OpenStudy (***[isuru]***):

u cant write tan (-945) as -tan(945) then u can write \[945 = 5\pi + \frac{ \pi }{ 4 }\] remember \[\pi = 180^{o}\] so... \[\tan(-945) = -\tan 945 = -\tan ( 5\pi + \frac{ \pi }{ 4 } ) = \tan [( 2\times 2\pi) + \pi + \frac{ \pi }{ 4 }]\] \[= - \tan ( \pi + \frac{ \pi }{ 4 } )\] \[= - \tan \frac{ \pi }{ 4 } = - \tan 45^{o} = -1\]

OpenStudy (***[isuru]***):

got it :) ?

OpenStudy (anonymous):

=tan(-270+45)=-tan(45)=-1

OpenStudy (anonymous):

thanks sir

OpenStudy (anonymous):

i got it

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

cos1degree.cos2degree.cos3degree...........cos179degree =??

OpenStudy (anonymous):

sir plz give me ans.

OpenStudy (anonymous):

it is <1

OpenStudy (anonymous):

\[ \tan( k \pi + x)= \tan(x), k=0,\pm 1\pm 2,\pm3 \cdots \\ \tan(-x)=-\tan(x)\\ \tan\left(- \left(\frac { \pi}4+5\pi \right) \right) =- \tan\left( \frac { \pi}4+5\pi \right) =- \tan\left( \frac { \pi}4 \right) =-1 \]